OpenStudy (anonymous):

Help please.. Evaluate the logarithm. log6 1/36

OpenStudy (anonymous):

\[\log_{6} \frac{ 1 }{ 36 }\]

OpenStudy (campbell_st):

write 1/36 as a power of 6 \[\frac{1}{36} = 6^{-2}\]

OpenStudy (campbell_st):

the base of the log and base of the exponent are the same so \[\log_{a} (a^x) = x\] apply the law to your question

OpenStudy (anonymous):